When the price of a product increases by 6% and the quantity demanded of the product changes from 1 400 to 1 200, what is the price elasticity of demand of the product? (Ignore the negative sign and round off to 2 decimal places)


1
Expert's answer
2021-09-10T12:20:39-0400

"\u0415 = \u2206Q \/ \u2206P"

∆Q(%) = (Q2-Q1)/Q1

∆P(%) = (P2-P1)/P1

∆P=6%

∆Q=(1200-1400)/1400=0.14=14%

"\u0415 = 14\/6=2.33"